# Spatio-Temporal Adaptive Diffusion Models for EEG Super-Resolution in Epilepsy Diagnosis

### v2

@@ Abstract
Electroencephalogram (EEG) technology, particularly high-density EEG (HD EEG) devices, is widely used in fields such as neuroscience. HD EEG devices improve the spatial resolution of EEG by placing more electrodes on the scalp, meeting the requirements of clinical diagnostic applications such as epilepsy focus localization. However, this technique faces challenges such as high acquisition costs and limited usage scenarios. In this paper, spatio-temporal adaptive diffusion models (STADMs) are proposed to pioneer the use of diffusion models for achieving spatial SR reconstruction from low-resolution (LR, 64 channels or fewer) EEG to high-resolution (HR, 256 channels) EEG. Specifically, a spatio-temporal condition module is designed to extract the spatio-temporal features of LR EEG, which then serve as conditional inputs to guide the reverse denoising process of diffusion models. Additionally, a multi-scale Transformer denoising module is constructed to leverage multi-scale convolution blocks and cross-attention-based diffusion Transformer blocks for conditional guidance to generate subject-adaptive SR EEG. Experimental results demonstrate that the proposed method effectively enhances the spatial resolution of LR EEG and quantitatively outperforms existing methods. Furthermore, STADMs demonstrate their value by applying synthetic SR EEG to classification and source localization tasks of epilepsy patients, indicating their potential to significantly improve the spatial resolution of LR EEG.
@@ 1 Introduction
Electroencephalogram (EEG) is a non-invasive, cost-effective neuroimaging technique with high temporal resolution, playing a crucial role in clinical diagnosis and cognitive neuroscience research. Unlike other neuroimaging techniques such as functional magnetic resonance imaging (fMRI) and positron emission tomography (PET), EEG offers superior temporal resolution and can continuously monitor brain activity in natural environments. This capability provides substantial support for studying dynamic cognitive processes [1]. In recent years, the emergence of portable and wireless EEG devices has further improved the flexibility and usability of EEG, enabling its widespread application in emerging neural engineering scenarios such as sleep monitoring and brain-computer interface (BCI) control [2, 3, 4]. However, these devices typically have fewer than 64 channels, resulting in lower spatial resolution of the acquired EEG data, which severely limits its potential in clinical diagnosis and brain function imaging research [5]. Currently, high-density (HD) EEG devices improve the spatial resolution of EEG by employing more electrode quantities, thus assisting doctors in performing evaluations or diagnoses in different clinical scenarios. For example, recent studies have shown that non-invasive HD EEG devices can objectively, non-invasively, and accurately determine the epileptic focus area [6, 7, 8]. However, HD EEG devices greatly increase hardware and acquisition costs, their bulkiness and complex wiring prevent subjects from moving their heads freely, restricting their use to controlled environments such as laboratories or hospitals. In addition, due to poor wearing comfort, subjects cannot be monitored for extended periods [9]. These factors collectively lead to difficulties in acquisition and limited application scenarios for HD EEG devices.
@@
The aforementioned observation highlights a significant dilemma in the current clinical diagnosis and treatment using EEG: due to high costs, discomfort from wearing, and other factors, high-resolution (HR) EEG is scarce. Consequently, the spatial resolution of low-resolution (LR) EEG is insufficient for clinical needs, particularly in the preoperative evaluation of patients with refractory epilepsy [10]. In such cases, the EEG spatial super-resolution (SR) method offers an effective remedy. This technique enables the reconstruction of HR EEG from LR EEG collected from subjects, thereby enhancing the spatial resolution of the existing EEG acquisition systems. For instance, EEG SR can reconstruct 256-channel EEG from 32-channel counterparts. Notably, EEG SR methods hold substantial potential for clinical applications, providing a cost-effective and efficient method to acquire HR EEG data.
@@
Currently, in the field of enhancing spatial resolution of EEG, the methodologies predominantly fall into two main categories: channel interpolation based on mathematical models, and EEG SR reconstruction utilizing deep learning methods. Channel interpolation methods principally rely on predefined scalp surface models to reconstruct missing channel data using signals from other channels, thereby improving spatial resolution. These methods involve calculating the second-order derivatives of the interpolation function in space to explore the distribution across the electrode channels [11]. Subsequent studies have attempted more precise measurements of the spatial distances between EEG electrodes and more accurate predictions for missing channel data, for example, using inverse distance weighting based on Euclidean distances [12] and multi-quadratic interpolation methods [13]. However, these methods only utilize the spatial correlations between adjacent electrodes and typically struggle to effectively capture and reconstruct the high-frequency and transient components of brain electrical activity, usually yielding limited results for non-stationary signals.
@@
In recent years, deep learning methods have become mainstream for analyzing and enhancing medical images [14, 15, 16]. Specifically, deep learning-based EEG SR technology involves training deep learning models to learn the implicit mappings between LR EEG and HR EEG, and using these trained models to reconstruct SR EEG. Notably, research has introduced EEG SR techniques based on convolution neural networks (CNN) [17], followed by studies proposing deep EEG SR frameworks using graph convolution networks (GCNs) [18] to correlate the structural and functional connections of participants’ brains [19]. Additionally, generative models offer a new learning paradigm for medical image reconstruction [20, 21]. Existing researches have utilized approaches such as Autoencoders (AE) [22] and Generative Adversarial Networks (GANs) [23] to enhance the spatial resolution of EEG. While existing deep learning-based EEG SR methods have achieved significant improvements in enhancing EEG spatial resolution, they are still limited in effectively capturing both the temporal dynamics and spatial features of EEG, resulting in limited generalizability and practicality of the models. Moreover, the enhanced spatial resolution achieved by these methods is constrained, with a maximum capability of reconstructing EEG data up to 64-channel level, which may not meet the requirements of specific clinical scenarios.
@@
In clinical diagnostic and therapeutic settings, achieving high-fidelity reconstruction of HR (256 channels) ground truth EEG presents a substantial challenge. The key is accurately depicting the dependencies between LR EEG and HR EEG and addressing the significant channel-level disparity between them. Recently, a novel generative model—diffusion models [24] have demonstrated formidable generative capabilities in the field of computer vision, sparking a surge in the Artificial Intelligence Generated Content (AIGC) [25, 26]. Moreover, diffusion models have been shown to surpass GANs in image generation quality, while also offering ideal characteristics such as distribution coverage, fixed training targets, and scalability [27]. Besides, Latent Diffusion Models (LDMs) [28] extend diffusion models into latent spaces, utilizing cross-attention mechanisms to achieve refined control in conditional generation tasks. Moreover, SynDiff [29] and DiffMDD [30] have both demonstrated that diffusion models can be applied to the generation of medical images, thereby improving the accuracy of disease diagnosis. Despite this, the application of diffusion models in enhancing the spatial resolution of EEG remains relatively unexplored. Therefore, this paper aims to investigate the potential of diffusion models in achieving EEG SR, particularly in enhancing spatial resolution. Through this novel generative model, we hope to provide a novel and effective method for processing and analyzing EEG data, supporting more accurate EEG studies and applications across various neuroscience contexts.
@@
In this paper, a novel EEG SR method, called spatio-temporal adaptive diffusion models (STADMs), is proposed to achieve spatial SR reconstruction from LR EEG to HR EEG. This method enhances the spatial resolution of EEG by learning the latent mapping relationship between LR EEG and HR EEG through diffusion learning strategy. The main contributions of this paper are summarized as follows:
@@
A spatio-temporal adaptive framework based on diffusion models is proposed for EEG SR reconstruction. To the best of our knowledge, this is the first instance where diffusion models are employed to achieve spatial SR reconstruction from LR EEG to HR EEG. This provides an efficient and cost-effective method for obtaining HR EEG, enhancing the portability and usability of HR EEG, and further advancing fields such as epileptic focus localization and BCI technology.
@@
The spatio-temporal condition (STC) module is designed to extract the spatio-temporal features of LR EEG. It further uses the generated spatio-temporal conditions to constrain the reverse process of the diffusion model, guiding the model to generate SR EEG that conforms to the characteristics of the subjects.
@@
The multi-scale Transformer denoising (MTD) module is constructed to map LR EEG to HR EEG. It extracts temporal features at various scales and uses cross-attention-based diffusion Transformer blocks to adaptively modulate the denoising process based on spatio-temporal conditions, addressing the channel-level disparity between LR EEG and HR EEG.
@@ 2.1 Preliminaries and Problem Statement
Before introducing our architecture, we briefly review some basic concepts needed to understand diffusion models [24]. Diffusion models assume a forward diffusion process which gradually applies noise to input data : , where are the variance schedule of the quantity of Gaussian noise.
In the reverse denoising process, the denoising model needs to learn the distributions to denoise the each state during training. Furthermore, the training of diffusion models is accomplished using a variational method-based training strategy [31]. Formally, the inference procedure can be described as a reverse Markov process that starts with Gaussian noise , and progresses towards the target data as:
Once is trained, new data can be sampled by initializing and sampling via the reparameterization trick, where denotes the number of forward process steps.
In this work, given an LR-HR EEG pair denoted as with being a degraded version of , where and denote the number of the LR and HR EEG channels, respectively. is the length of each EEG epoch. Therefore, in EEG SR tasks, the end-to-end mapping from the LR EEG to the corresponding SR EEG can be expressed as
where represent synthetic SR EEG. denotes the mapping function and represent the optimal parameters that are looked for to reconstruct the HR EEG.
@@ 2.2 Basic ideas
Learning the mapping relationship from LR EEG to HR EEG is challenging. To ensure that the generated results conform to the subject’s characteristics, the proposed STADMs use the spatio-temporal features of LR EEG as conditions and employs a cross-attention-based conditional injection mechanism to guide the denoising model for progressive denoising restoration. This mechanism embeds the spatio-temporal information of LR EEG into the high-dimensional latent space, indirectly guiding the model to learn the complex mapping relationship between LR EEG and HR EEG. This ensures that the synthetic SR EEG is highly consistent with HR EEG in both temporal and spatial domains, addressing the issue of the large channel-level disparity between LR EEG and HR EEG.
As Fig. 1 shows, the entire model framework consists of EEG pre-trained autoencoder, STC module and MTD module. During the training phase of STADMs, the HR EEG ground truth is first input into the EEG pre-trained encoder to obtain the corresponding latent vectors , and the state of each time step is gradually obtained through the forward process of the diffusion models. In the reverse process, the STC module first encodes the time series and channel spatial distribution matrix from LR EEG into the spatio-temporal representation conforming to the subject, and concatenates it with time embedding as the conditioning vector. MTD module takes the conditioning vectors and noise sampled from the standard normal distribution as input, implements the conditional guidance mechanism through the cross-attention layer, and predicts the noise at each step of the reverse process. The MTD module is trained by optimizing the mean squared error loss between the predicted noise and the actual noise, and finally generates SR EEG through progressive denoising. During the sampling phase of STADMs, the deterministic MTD module precisely generates SR EEG matching the input LR EEG by progressively sampling using the spatio-temporal conditional information of LR EEG and noise sampled from the Gaussian distribution.
@@ 2.3.1 Pre-trained EEG Autoencoders
LDMs [28] have demonstrated that mapping input data into a latent space through a pre-trained encoder can significantly enhance the quality of high-resolution natural image generation and improve performance on other downstream tasks. Following the work of [32], we employ a Masked Autoencoder (MAE) for asymmetric latent space representation of HR EEG. Given that the input EEG consists of multi-channel time series signals, we utilize two-dimensional EEG data as input, where each row represents the time series data of a single channel, and each column represents the values of all channels at a specific time point. Initially, this module divides the EEG data of each channel into fixed-length windows and randomly masks them according to a specified percentage. The MAE then predicts the missing values based on the surrounding contextual cues. By reconstructing the masked signals, the pre-trained EEG encoder can learn the latent spatio-temporal representations of brain activity across different subjects. Detailed information on the architecture of the pre-trained EEG encoder can be found in [32].
@@ 2.3.2 Spatio-temporal Condition Module
In EEG signals, each channel represents the signals collected by electrodes placed at specific locations on the brain, recording electrical activity from different brain regions [33, 34]. The spatio-temporal condition module based on Transformer is constructed to capture the temporal correlations between time points and the spatial correlations between EEG channels. This module takes the time series from LR EEG and the spatial positions of the channels as inputs and outputs corresponding encoding vectors as conditional information for the denoising model, where represents the LR EEG. This module can extract the spatio-temporal features of LR EEG, providing richer and more precise contextual information for the denoising model.
As the yellow part of Fig. 1 shows, the STC module is mainly composed of a spatial position embedding layer, a 1D convolution block, and a Transformer block. More specifically, the channel spatial embedding layer encodes the channel spatial coordinates of LR EEG, outputting a set of spatial features. Then, we partition the input time series into different patches and feed them into the 1D convolutional block to preliminarily extract temporal features [35]. Finally, the Transformer block concatenates the temporal features with position encoding and channel spatial embeddings to output the final spatio-temporal condition . The 1D convolution block consists of a 1D convolution layer, a ReLU activation function, and batch normalization layer. The Transformer block consists of a multi-head self-attention (MSA) layer followed by feed-forward networks, both wrapped in residual connections and layer normalization.

### v3

@@ Abstract
Electroencephalogram (EEG) technology, particularly high-density EEG (HD EEG) devices, are widely used in fields such as neuroscience. HD EEG devices improve the spatial resolution of EEG by placing more electrodes on the scalp, which meet the requirements of clinical diagnostic applications such as epilepsy focus localization. However, this technique faces challenges, such as high acquisition costs and limited usage scenarios. In this paper, spatio-temporal adaptive diffusion models (STADMs) are proposed to pioneer the use of diffusion models for achieving spatial SR reconstruction from low-resolution (LR, 64 channels or fewer) EEG to high-resolution (HR, 256 channels) EEG. Specifically, a spatio-temporal condition module is designed to extract the spatio-temporal features of LR EEG, which then used as conditional inputs to direct the reverse denoising process. Additionally, a multi-scale Transformer denoising module is constructed to leverage multi-scale convolution blocks and cross-attention-based diffusion Transformer blocks for conditional guidance to generate subject-adaptive SR EEG. Experimental results demonstrate that the STADMs significantly enhances the spatial resolution of LR EEG and quantitatively outperforms existing methods. Furthermore, STADMs demonstrate their value by applying synthetic SR EEG to classification and source localization tasks for epilepsy patients, indicating their potential to Substantially boost the spatial resolution of EEG.
@@ 1 Introduction
Electroencephalogram (EEG) is a economical, non-invasive neuroimaging technology with high temporal resolution, playing a vital role in cognitive neuroscience research and clinical diagnosis. Unlike other neuroimaging techniques such as positron emission tomography (PET) and functional magnetic resonance imaging (fMRI), EEG offers superior temporal resolution and can continuously monitor brain activity in natural environments. This capability substantially supports the study of dynamic cognitive processes [1]. Recently, the development of portable and wireless EEG devices has further improved the flexibility and usability of EEG, enabling its widespread application in emerging neural engineering scenarios such as sleep monitoring and brain-computer interface (BCI) control [2, 3, 4]. However, these devices typically have fewer than 64 channels, resulting in lower spatial resolution of the acquired EEG data, which severely limits its potential in clinical diagnosis and brain function imaging research [5]. Currently, high-density (HD) EEG devices improve the spatial resolution of EEG by employing more electrode quantities, thus assisting doctors in performing evaluations or diagnoses in different clinical scenarios. For example, recent studies have shown that non-invasive HD EEG devices can objectively, non-invasively, and accurately determine the epileptic focus area [6, 7, 8]. However, HD EEG devices greatly increase hardware and acquisition costs, their bulkiness and complex wiring prevent subjects from moving their heads freely, restricting their use to controlled environments such as laboratories or hospitals. In addition, due to poor wearing comfort, subjects cannot be monitored for extended periods [9]. These factors collectively lead to difficulties in acquisition and limited application scenarios for HD EEG devices.
@@
The aforementioned observation highlights a significant dilemma in the current clinical diagnosis and treatment using EEG: due to high costs, discomfort from wearing, and other factors, high-resolution (HR) EEG is scarce. Consequently, low-resolution (LR) EEG is insufficient for clinical needs, particularly in the pre-surgery assessment of patients with refractory epilepsy [10]. In such cases, the EEG spatial super-resolution (SR) method offers an effective remedy. This technique enables the reconstruction of HR EEG from LR EEG collected from subjects, thereby enhancing the spatial resolution of the existing EEG acquisition systems. For instance, EEG SR can reconstruct 256-channel EEG from 32-channel counterparts. Notably, EEG SR methods hold substantial potential for clinical applications, providing a cost-effective and efficient method to acquire HR EEG data.
@@
Currently, in the field of enhancing spatial resolution of EEG, the methodologies predominantly fall into two main categories: channel interpolation based on mathematical models and EEG SR reconstruction utilizing deep learning methods. Channel interpolation methods principally rely on predefined scalp surface models to reconstruct missing channel data using signals from other channels, thereby improving spatial resolution. These methods involve calculating the second-order derivatives of the interpolation function in space to explore the distribution across the electrode channels [11]. Subsequent studies have attempted more precise measurements of the spatial distances between EEG electrodes and more accurate predictions for missing channel data, for example, using inverse distance weighting based on Euclidean distances [12] and multi-quadratic interpolation methods [13]. However, these methods only utilize the spatial correlations among adjacent electrodes and typically struggle to effectively capture and reconstruct the high-frequency and transient components of brain electrical activity, resulting in suboptimal reconstruction performance for non-stationary EEG signals.
@@
Lately, deep learning approaches have become mainstream for analyzing and enhancing medical images [14, 15, 16, 17]. Specifically, deep learning-based EEG SR technology involves training deep learning models to learn the implicit mappings between LR EEG and HR EEG, and using these trained models to reconstruct SR EEG. Notably, research has introduced EEG SR techniques based on convolution neural networks (CNN) [18], followed by studies proposing deep EEG SR frameworks using graph convolution networks (GCNs) [19, 20] to correlate the structural and functional connections of participants’ brains [21]. Additionally, generative models offer a new learning paradigm for medical image reconstruction [22, 23, 24], such as Autoencoders (AE) and Generative Adversarial Networks (GANs) [25, 26, 27]. Existing researches have utilized approaches such as AE and GANs to enhance the spatial resolution of EEG [28, 29]. Although current EEG SR methods based on deep learning have made considerable advancements in enhancing EEG spatial resolution, they are still limited in effectively capturing both the temporal dynamics and spatial features of EEG, resulting in limited generalizability and practicality of the models. Moreover, the enhanced spatial resolution achieved by these methods is constrained, with a maximum capability of reconstructing EEG data up to 64-channel level, which may not meet the requirements of specific clinical scenarios.
@@
In clinical diagnostic and therapeutic settings, achieving high-fidelity reconstruction of HR (256 channels) ground truth EEG presents a substantial challenge. The key is accurately depicting the dependencies between LR EEG and HR EEG and addressing the significant channel-level disparity between them. Recently, a novel generative model—diffusion models [30] have exhibited remarkable generative capabilities in computer vision, leading to a significant increase in the Artificial Intelligence Generated Content (AIGC) [31, 32, 33]. Moreover, diffusion models have demonstrated superior image generation quality compared to GANs, while also providing advantageous features such as comprehensive distribution coverage, extensibility, and stable training targets [34]. Besides, Latent Diffusion Models (LDMs) [35] extend diffusion models into latent spaces, utilizing cross-attention mechanisms to achieve refined control in conditional generation tasks. Furthermore, SynDiff [36] , Brain Diffuser [37] and DiffMDD [38] have illustrated that diffusion models are applicable in generating medical images, thereby advancing the accuracy of disease diagnosis. Despite this, Applying diffusion models to enhance EEG spatial resolution has not been thoroughly explored. Therefore, this paper aims to investigate the potential of diffusion models in achieving EEG SR, particularly in enhancing spatial resolution. Through this novel generative model, we hope to provide a novel and effective method for processing and analyzing EEG data, supporting more accurate EEG studies and applications across various neuroscience contexts.
@@
In this paper, a novel EEG SR method, called spatio-temporal adaptive diffusion models (STADMs), is proposed to achieve spatial SR reconstruction from LR EEG to HR EEG. Employing a diffusion learning strategy, This approach improves EEG spatial resolution by learning the latent mapping relationship between LR EEG and HR EEG. The primary contributions of this work are summarized as:
@@
A spatio-temporal adaptive framework based on diffusion models is proposed for EEG SR reconstruction. According to our knowledge, this marks the first time diffusion models are employed to achieve spatial SR reconstruction from LR EEG to HR EEG. This provides an efficient and cost-effective method for obtaining HR EEG, enhancing the portability and usability of HR EEG, and further advancing fields such as epileptic focus localization and BCI technology.
@@
The spatio-temporal condition module (STCM) is designed to capture the spatio-temporal features of LR EEG. It further uses the generated spatio-temporal conditions to constrain the reverse denoising process, guiding the model to generate SR EEG that conforms to the characteristics of the subjects.
@@
The multi-scale Transformer denoising module(MTDM) is constructed to map LR EEG to HR EEG. It extracts temporal features at various scales and uses cross-attention-based diffusion Transformer blocks to adaptively modulate the denoising process based on spatio-temporal conditions, addressing the channel-level disparity between LR EEG and HR EEG.
@@ 2.1 Preliminaries and Problem Statement
We will provide a brief review of some fundamental concepts essential for understanding diffusion models [30] before introducing our architecture. Diffusion models presuppose a forward diffusion process that incrementally adds noise to input data as , where signifies the variance schedule for the amount of Gaussian noise.
In the reverse denoising process, the denoising model needs to learn the distributions to denoise the each state during training. Furthermore, the training of diffusion models is accomplished using a variational method-based training strategy [39]. Formally, the inference procedure can be characterized as a reverse Markov process that starts with Gaussian noise , and progresses towards the target data as:
With trained, new target data can be obtained by initializing the variable and then using the reparameterization trick to sample step by step, where denotes the number of forward process steps.
In this work, given an LR-HR EEG pair denoted as with being a degraded version of , where and denote the number of the LR and HR EEG channels, respectively. is the length of each EEG epoch. Therefore, the process of mapping the LR EEG to the corresponding SR EEG can be described as
where represents synthetic SR EEG. stands for the mapping function, while represents the optimal parameters sought to reconstruct the HR EEG.
@@ 2.2 Basic ideas
Learning the mapping relationship from LR EEG to HR EEG is challenging. To ensure that the generated results conform to the subject’s characteristics, the proposed STADMs use the spatio-temporal features of LR EEG as conditions and employs a cross-attention-based conditional injection mechanism to guide the denoising model for progressive denoising restoration. This mechanism embeds the spatio-temporal information of LR EEG into the high-dimensional latent space, indirectly guiding the model to comprehend the potential correlation between LR EEG and HR EEG. This ensures that the synthetic SR EEG is highly consistent with HR EEG in both temporal and spatial domains, addressing the issue of the large channel-level disparity between LR EEG and HR EEG.
As Fig. 1 shows, the entire model framework consists of EEG pre-trained autoencoder, STCM and MTDM. During the training phase of STADMs, the HR EEG ground truth is first input into the EEG pre-trained encoder to obtain the corresponding latent vectors . The forward diffusion process progressively derives the state of each time step . In the reverse denoising process, the STCM first encodes the time series and channel spatial distribution matrix from LR EEG into the spatio-temporal representation conforming to the subject, and concatenates it with time embedding as the conditioning vector. MTDM takes the conditioning vectors and noise sampled from the standard normal distribution as input, implements the conditional guidance mechanism through the cross-attention layer, and predicts the noise at each step of the reverse process. To train the MTDM, the Mean Squared Error (MSE) loss is minimized between the predicted and actual noise at each step, ultimately generating SR EEG through progressive denoising. During the sampling phase of STADMs, the deterministic MTDM precisely generates SR EEG matching the input LR EEG by progressively sampling using the spatio-temporal conditional information of LR EEG and noise sampled from the Gaussian distribution.
@@ 2.3.1 Pre-trained EEG Autoencoders
LDMs [35] have demonstrated that mapping input data into a latent space through a pre-trained encoder can significantly enhance the quality of high-resolution natural image generation and improve performance on other downstream tasks. Following the work of [40], we employ a Masked Autoencoder (MAE) for asymmetric latent space representation of HR EEG. Given that the input EEG consists of multi-channel time series signals, we utilize two-dimensional EEG data as input, where each row represents the time series data of a single channel, and each column represents the values of all channels at a specific time point. Initially, this module divides the EEG data of each channel into fixed-length windows and randomly masks them according to a specified percentage. The MAE then predicts the missing values using the surrounding contextual cues. By reconstructing the masked signals, the pre-trained EEG encoder can learn the latent spatio-temporal representations of brain activity across different subjects. Detailed information on the architecture of the pre-trained EEG encoder can be found in [40].
@@ 2.3.2 Spatio-temporal Condition Module
In EEG signals, each channel represents the signals collected by electrodes placed at specific locations on the brain, recording electrical activity from different brain regions [41, 42]. The spatio-temporal condition module based on Transformer is constructed to capture the temporal correlations between time points and the spatial correlations between EEG channels. This module takes the time series from LR EEG and the spatial positions of the channels as inputs and outputs corresponding encoding vectors as conditional information for the reverse denoising process, where represents the LR EEG. By extracting the spatio-temporal features of LR EEG, this module provides richer and more precise contextual information for the denoising model.
As the yellow part of Fig. 1 shows, the STCM is mainly composed of a spatial position embedding layer, a 1D convolution block, and a Transformer block. More specifically, the channel spatial embedding layer encodes the channel spatial coordinates of LR EEG, outputting a set of spatial features. Then, we partition the input time series into different patches and process them with the 1D convolution block for the initial extraction of temporal features [43]. Finally, the Transformer block concatenates the temporal features with position encoding and channel spatial embeddings to output the final spatio-temporal condition . The 1D convolution block consists of a 1D convolution layer, batch normalization layer and a ReLU activation function. In the Transformer block, a multi-head self-attention (MSA) layer is followed by feed-forward networks, both of which are encapsulated with residual connections and layer normalization.